# Nonstandard Universes

### v1

@@ 6.1 Existence of Nonstandard Frameworks
Let $V(X)$ be a superstructure over an infinite base set $X$. Then there exists a transfer map (i.e. a bounded elementary embedding) $V(X)\buildrel*\over{\to}V(Y)$ of $V(X)$ into some superstructure $V(Y)$ with base set $Y$ such that
@@ 6.1.1 Ultrapower Proof
We start from a superstructure $V(X)$ over a base set $X$. Let $\mathcal{U}$ be a countably incomplete ultrafilter over some index set $I$. The construction of the nonstandard framework proceeds over 9 steps. The first 8 steps will define a transfer map $*:V(X)\to V(Y)$ from the superstructure $V(X)$ to a superstructure $V(Y)$ over base set $Y$. These steps do not require $\mathcal{U}$ to be countably incomplete. In the $9^{\text{th}}$ step, the countable incompleteness of $\mathcal{U}$ is used to show that * induces a nonstandard framework.
@@
i.e. $Y$ is just the ultrapower $X^{I}/\mathcal{U}$. We shall assume that $I$ is chosen so that $Y$ is a base set for $V(Y)$. This can be done as follows: Suppose that $X$ is a set of rank $\beta$, and let $I$ be a set of rank $\gamma\geq\beta+\omega$. Let $\mathcal{U}$ be an ultrafilter over $I$, and let $f:I\to X$. If $\gamma$ is a successor ordinal, $\gamma=\delta+1$, then $I$ has an element $i_{0}$ of rank $\delta$, but no elements of higher rank. It follows easily that the rank of $f$ is $\delta+3=\gamma+2$. On the other hand, if $\gamma$ is a limit ordinal, then $\sup_{i\in I}\text{rank}(i)=\gamma$, but the supremum is not attained. Hence $f$ has rank $\gamma$. It therefore follows that every $f:I\to X$ has the same rank, namely either $\gamma+2$ if $\gamma$ is a successor ordinal, or $\gamma$ if it is limit. Now an element of an element of $Y=X^{I}/\mathcal{U}$ is precisely a function $f:I\to X$. Thus all elements of elements of $Y$ have the same rank. The argument in Example 1.6 now shows that $Y$ is a base set, i.e. that $y\cap V(Y)=\varnothing$ for all $y\in Y$. Step 2: Now we define a structure $(W,=_{\mathcal{U}},\in_{\mathcal{U}})$ which forms part of an intermediate step in the definition of the transfer map $*:V(X)\to V(Y)$: Define binary relations $=_{\mathcal{U}}$ and $\in_{\mathcal{U}}$ on $V(X)^{I}$ as follows:
@@
    $g=_{\mathcal{U}}f$ if and only if $g/\mathcal{U}=f/\mathcal{U}$
@@
To decompose this definition, define $\iota:V(X)\to W:a\mapsto c_{a}$. Then if $a\in V_{n}(X)$, $\iota(a)=c_{a}\in W_{n}$. Then $*:V(X)\to V(Y)$ is just the composition
[…]
@@
We have now proved ()($\ddagger$)( ‡ ). In particular, we have
@@
But since $c_{a}(i)=a$ for all $i\in I$, the set $\{i\in I:V(X)\vDash\varphi[c_{a_{1}}(i),\dots,c_{a_{n}}(i)]\}$ is either all of $I$, in which case $V(X)\vDash\varphi[a_{1},\dots,a_{n}]$, or it is $\varnothing$, in which case $V(X)\vDash\lnot\varphi[a_{1},\dots,a_{n}]$. This proves that ()($\dagger$)( † ), i.e. that $\cdot/\mathcal{U}$ is a transfer map for the language $\mathcal{L}_{\in}$.
Step 9: In order to complete the proof of Theorem 6.1, it remains to show that $*:V(X)\to V(Y)$ is a nonstandard framework over $X$, i.e. that there is a countable subset $A$ of $X$ such that ${}^{\sigma}A:=\{{}^{*}a:a\in A\}$ is a proper subset of ${}^{*}A$. This is the only place where we need the fact that the ultrafilter $\mathcal{U}$ is countably incomplete. In fact, we can directly prove:
@@ 6.1.2 Compactness Theorem Proof
We are now able to apply Theorem 6.5 to deduce that there is a bounded elementary extension $h:\mathfrak{A}\hookrightarrow(V(Y),\in)$. Let $*:=h\restriction V(X)$. Then $*:(V(X),\in)\hookrightarrow(V(Y),\in)$ is a composition of two bounded elementary embeddings, and thus a bounded elementary embedding, with ${}^{*}x=x$ for all $x\in X$, and ${}^{*}X=Y$. Finally, if $U\subseteq X$ is infinite, there is $b\in B-V(X)$ such that $bEU$. As $V(X)\vDash\forall u\in c_{U}\;(u\in c_{X})[U,X]$, we also have that $bEU$ implies $bEX$ for all $b\in B$, and hence $bEU$ implies $b\in A$, by definition of truncation. Thus $h(b)$ is defined, and $h(b)\in h(U)={}^{*}U$. The fact that $b\not\in X$ means that $h(b)\not\in\{h(x):x\in X\}=\{^{*}x:x\in X\}$, so $\{^{*}x:x\in U\}$ is a proper subset of ${}^{*}U$.
@@ 6.2 Existence of Enlargements
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement. Then $|\mathbb{N}^{*}|\geq\sup\{|A|:A\in\mathbb{U},|A|<\kappa\}$.
Proof: Suppose that $A\in\mathbb{U}$ has cardinality $<\kappa$. By Theorem 5.10, there is a hyperfinite set $B$ such that ${}^{\sigma}A\subseteq B\subseteq{}^{*}A$. As $B$ is hyperfinite, there is a $n\in{}^{*}\mathbb{N}$ and a bijection $f:\{0,1,\dots,n-1\}\to B$, and thus an injection $h:B\hookrightarrow{}^{*}\mathbb{N}$. Thus ${}^{*}\mathbb{N}$ has a subset of cardinality $|B|$, and thus one of cardinality $|A|=|{}^{\sigma}A|\leq|B|$. Hence $|{}^{*}\mathbb{N}|\geq|A|$ for any $A\in\mathbb{U}$ with $|A|<\kappa$.
@@
Using the above lemma, it can be seen that the ultrapower construction does not automatically provide enlargements. For example, consider the ultrapower construction $*:V(X)\to V(Y)$ over an infinite base set $X$, with $Y=X^{I}/\sim_{\mathcal{U}}$ and $I=\mathbb{N}$. we may assume that $\mathbb{N}\subseteq X$. Then ${}^{*}\mathbb{N}$ is of the form $c_{\mathbb{N}}/\mathcal{U}=\{g/\mathcal{U}:g\in_{\mathcal{U}}\mathbb{N}\}$, and so $|{}^{*}\mathbb{N}|\leq|\mathbb{N}^{\mathbb{N}}|=2^{\aleph_{0}}$. Now $\mathbb{N},\mathcal{P}(\mathbb{N}),\mathcal{P}\mathcal{P}(\mathbb{N})\dots$ are all members of $\mathbb{U}$, so if $\kappa>(2^{\aleph_{0}})^{+}$, then $\sup\{|A|:A\in\mathbb{U},|A|<\kappa\}>2^{\aleph_{0}}$. Hence if $\kappa>(2^{\aleph_{0}})^{+}$, then * cannot be a $\kappa$-enlargement.
@@
[…]
@@
and hence $I_{\{B\}}\subseteq\{a\in I:f(a)\in B\}$. It follows that $f/\mathcal{U}\in c_{B}/\mathcal{U}={}^{*}B$. As $B\in\mathcal{B}$ is arbitrary, it follows that $f/\mathcal{U}\in\bigcap_{B\in\mathcal{B}}{}^{*}B=\bigcap{}^{\sigma}\mathcal{B}$. As $\mathcal{B}$ is arbitrary, it follows that $\bigcap{}^{\sigma}\mathcal{B}\neq\varnothing$ for all $\mathcal{B}\in V(X)$ with the f.i.p, and thus by Theorem 5.3, the extension $*:V(X)\to V(Y)$ is an enlargement.
@@ 6.3 Existence of Comprehensive Transfer Maps
Let $V(X)$ be a superstructure over $X$, and let $*:V(X)\to V({}^{*}X)$ be the transfer map provided by an ultrapower construction, as in Section 6.1.1. Then * is comprehensive.
@@ 6.4 Existence of Polysaturated Extensions via Ultrapowers
We first show that the following claim holds: Claim: If $\mathcal{L}$ is a countable language, and if $\Sigma(x)$ is a set of $\mathcal{L}$-formulas which is finitely satisfiable in an ultraproduct $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$, then $\Sigma(x)$ is satisfiable. So suppose that $\Sigma(x)$ is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Since $\mathcal{L}$ is countable, so is $\Sigma(x)$, and hence we can enumerate it:
@@
Suppose now that we have a family$\{\mathfrak{A}_{i}:i\in I\}$ of $\mathcal{L}$-structures, indexed by a set $I$, and an ultrafilter $\mathcal{U}$ on $I$. We seek conditions on $\mathcal{U}$ which will ensure that the ultraproduct $\prod_{I}A_{i}/\mathcal{U}$ is $\kappa$-saturated for $\kappa\geq\omega_{1}$. Moreover, as in Theorem 6.10, we seek a condition on $\mathcal{U}$, i.e. one which is independent of the models $\mathfrak{A}_{i}$.
@@
Observe that if $\mathcal{U}$ satisfies (S) then any ultraproduct modulo $\mathcal{U}$ interpreting a language $\mathcal{L}$ of cardinality $<\kappa$ is $\kappa$-saturated. To see this, suppose that $X\subseteq\prod_{I}A_{i}/\mathcal{U}$ has $|X|<\kappa$. Then $\mathcal{L}_{X}$ is also a language of cardinality $<\kappa$. Automatically, therefore, any set of $\mathcal{L}_{X}$-formulas $\Sigma(x)$ has $|\Sigma(x)|<\kappa$. Condition (S) then immediately yields that any set of $\mathcal{L}_{X}$-formulas $\Sigma(x)$ which is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$.
@@
Thus let $\Sigma(x)$ be a set of formulas of cardinality $<\kappa$ which is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Then if $\Theta$ is a finite subset of $\Sigma$, the set $\{i\in I:\mathfrak{A}_{i}\vDash\exists x(\bigwedge\Theta)\}$ belongs to $\mathcal{U}$, by Łos’ Theorem. We thus have a map
@@
We say that a sequence $(\Phi_{i})_{i\in I}$ of members of $\mathcal{P}^{<\omega}(\Sigma)$ supports $p$ if it satisfies (i), (ii) of Lemma 6.11. Thus if there is a sequence supporting the relation $p$ defined in ()($\star$)( ⋆ ) above, then $\Sigma$ is satisfiable in $\prod_{i\in I}\mathfrak{A}_{i}/\mathcal{U}$.
@@
Observe that the order-reversal defined in ()($\star$)( ⋆ ) need not be either anti-additive or locally finite. However, if an abstract order-reversal $p$ has a support $\Phi:=(\Phi_{i})_{i\in I}$, then we can define another order-reversal $p_{\Phi}$ which has those properties:
@@
We now seek conditions that will ensure that any order-reversal on any set $X$ of cardinality $<\kappa$ has a support. Observe that if $p,p^{\prime}$ are order-reversals so that $p^{\prime}\leq p$, then $p^{\prime}$ is locally finite if $p$ is, since $\{\Theta:i\in p^{\prime}(\Theta)\}\subseteq\{\Theta:i\in p(\Theta)\}$ for all $i\in I$. Furthermore, if $\Phi$ is a support of $p^{\prime}$, then it is a support of $p$ also, since $I=\{i:i\in p^{\prime}(\Phi_{i})\}\subseteq\{i:i\in p(\Phi_{i})\}$. Hence if we can define for every order-reversal $p$ two smaller order -eversals $Lp\leq p,Cp\leq p$ so that $Lp$ is locally finite, and $Cp$ is anti-additive, then $CLp\leq Lp\leq p$ is anti-additive and locally finite. Hence $CLp$ has a support, and this will be a support of $p$ also.
@@
  • Suppose that $\mathcal{U}$ is a countably incomplete $\kappa$-good ultrafilter on a set $I$, and that $(\mathfrak{A}_{i})_{i\in I}$ is a family of models interpreting a language $\mathcal{L}$ of cardinality $<\kappa$.
@@
If $\mathcal{U}$ is a countably incomplete $\kappa$-good ultrafilter, then every ultraproduct modulo $\mathcal{U}$ interpreting a language of cardinality $<\kappa$ is $\kappa$-saturated.

### v2

@@ 6.1 Existence of Nonstandard Frameworks
Let $V(X)$ be a superstructure over an infinite base set $X$. Then there exists a transfer map (i.e. a bounded elementary embedding) $V(X)\mathrel{\mathop{\kern 0.0pt\to}\limits^{*}}V(Y)$ of $V(X)$ into some superstructure $V(Y)$ with base set $Y$ such that
@@ 6.1.1 Ultrapower Proof
We start from a superstructure $V(X)$ over a base set $X$. Let $\mathcal{U}$ be a countably incomplete ultrafilter over some index set $I$. The construction of the nonstandard framework proceeds over 9 steps. The first 8 steps will define a transfer map $V(X)\to V(Y)$ from the superstructure $V(X)$ to a superstructure $V(Y)$ over base set $Y$. These steps do not require $\mathcal{U}$ to be countably incomplete. In the $9^{\text{th}}$ step, the countable incompleteness of $\mathcal{U}$ is used to show that * induces a nonstandard framework.
@@
i.e. $Y$ is just the ultrapower $X^{I}/\mathcal{U}$. We shall assume that $I$ is chosen so that $Y$ is a base set for $V(Y)$. This can be done as follows: Suppose that $X$ is a set of rank $\beta$, and let $I$ be a set of rank $\gamma\geq\beta+\omega$. Let $\mathcal{U}$ be an ultrafilter over $I$, and let $f:I\to X$. If $\gamma$ is a successor ordinal, $\gamma=\delta+1$, then $I$ has an element $i_{0}$ of rank $\delta$, but no elements of higher rank. It follows easily that the rank of $f$ is $\delta+3=\gamma+2$. On the other hand, if $\gamma$ is a limit ordinal, then $\sup_{i\in I}\text{rank}(i)=\gamma$, but the supremum is not attained. Hence $f$ has rank $\gamma$. It therefore follows that every $f:I\to X$ has the same rank, namely either $\gamma+2$ if $\gamma$ is a successor ordinal, or $\gamma$ if it is limit. Now an element of an element of $Y=X^{I}/\mathcal{U}$ is precisely a function $f:I\to X$. Thus all elements of elements of $Y$ have the same rank. The argument in Example 1.6 now shows that $Y$ is a base set, i.e. that $y\cap V(Y)=\varnothing$ for all $y\in Y$.
Step 2: Now we define a structure $(W,=_{\mathcal{U}},\in_{\mathcal{U}})$ which forms part of an intermediate step in the definition of the transfer map $V(X)\to V(Y)$: Define binary relations $_{\mathcal{U}}$ and $\in_{\mathcal{U}}$ on $V(X)^{I}$ as follows:
@@
    $g=_{\mathcal{U}}f$ if and only if $g/\mathcal{U}=f/\mathcal{U}$ .
@@
To decompose this definition, define $\iota:V(X)\to W:a\mapsto c_{a}$. Then if $a\in V_{n}(X)$, $\iota(a)=c_{a}\in W_{n}$. Then $V(X)\to V(Y)$ is just the composition
$V(X)\overset{\iota}{\hookrightarrow}W\overset{\cdot/\mathcal{U}}{\rightarrow}V(Y).$
@@
We have now proved $(\ddagger)$. In particular, we have
@@
But since $c_{a}(i)=a$ for all $i\in I$, the set $\{i\in I:V(X)\vDash\varphi[c_{a_{1}}(i),\dots,c_{a_{n}}(i)]\}$ is either all of $I$, in which case $V(X)\vDash\varphi[a_{1},\dots,a_{n}]$, or it is $\varnothing$, in which case $V(X)\vDash\lnot\varphi[a_{1},\dots,a_{n}]$. This proves that $(\dagger)$, i.e. that $\cdot/\mathcal{U}$ is a transfer map for the language $\mathcal{L}_{\in}$.
Step 9: In order to complete the proof of Theorem 6.1, it remains to show that $V(X)\to V(Y)$ is a nonstandard framework over $X$, i.e. that there is a countable subset $A$ of $X$ such that $^{\sigma}A:=\{{}^{*}a:a\in A\}$ is a proper subset of $^{*}A$. This is the only place where we need the fact that the ultrafilter $\mathcal{U}$ is countably incomplete. In fact, we can directly prove:
@@ 6.1.2 Compactness Theorem Proof
We are now able to apply Theorem 6.5 to deduce that there is a bounded elementary extension $h:\mathfrak{A}\hookrightarrow(V(Y),\in)$. Let $h\restriction V(X)$. Then $(V(X),\in)\hookrightarrow(V(Y),\in)$ is a composition of two bounded elementary embeddings, and thus a bounded elementary embedding, with $^{*}x=x$ for all $x\in X$, and $^{*}X=Y$. Finally, if $U\subseteq X$ is infinite, there is $b\in B-V(X)$ such that $bEU$. As $V(X)\vDash\forall u\in c_{U}\;(u\in c_{X})[U,X]$, we also have that $bEU$ implies $bEX$ for all $b\in B$, and hence $bEU$ implies $b\in A$, by definition of truncation. Thus $h(b)$ is defined, and $h(b)\in h(U)={}^{*}U$. The fact that $b\not\in X$ means that $h(b)\not\in\{h(x):x\in X\}=\{^{*}x:x\in X\}$, so $\{^{*}x:x\in U\}$ is a proper subset of $^{*}U$.
@@ 6.2 Existence of Enlargements
Suppose that $\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement. Then $|\mathbb{N}^{*}|\geq\sup\{|A|:A\in\mathbb{U},|A|<\kappa\}$.
Proof: Suppose that $A\in\mathbb{U}$ has cardinality $\kappa$. By Theorem 5.10, there is a hyperfinite set $B$ such that $^{\sigma}A\subseteq B\subseteq{}^{*}A$. As $B$ is hyperfinite, there is a $n\in{}^{*}\mathbb{N}$ and a bijection $f:\{0,1,\dots,n-1\}\to B$, and thus an injection $h:B\hookrightarrow{}^{*}\mathbb{N}$. Thus $^{*}\mathbb{N}$ has a subset of cardinality $|B|$, and thus one of cardinality $|A|=|{}^{\sigma}A|\leq|B|$. Hence $|{}^{*}\mathbb{N}|\geq|A|$ for any $A\in\mathbb{U}$ with $|A|<\kappa$.
@@
Using the above lemma, it can be seen that the ultrapower construction does not automatically provide enlargements. For example, consider the ultrapower construction $V(X)\to V(Y)$ over an infinite base set $X$, with $Y=X^{I}/\sim_{\mathcal{U}}$ and $I=\mathbb{N}$. we may assume that $\mathbb{N}\subseteq X$. Then $^{*}\mathbb{N}$ is of the form $c_{\mathbb{N}}/\mathcal{U}=\{g/\mathcal{U}:g\in_{\mathcal{U}}\mathbb{N}\}$, and so $|{}^{*}\mathbb{N}|\leq|\mathbb{N}^{\mathbb{N}}|=2^{\aleph_{0}}$. Now $\mathbb{N},\mathcal{P}(\mathbb{N}),\mathcal{P}\mathcal{P}(\mathbb{N})\dots$ are all members of $\mathbb{U}$, so if $\kappa>(2^{\aleph_{0}})^{+}$, then $\sup\{|A|:A\in\mathbb{U},|A|<\kappa\}>2^{\aleph_{0}}$. Hence if $\kappa>(2^{\aleph_{0}})^{+}$, then * cannot be a $\kappa$-enlargement.
@@
Given a base set $X$, let $I:=\mathcal{P}^{<\omega}(V(X))$ be the family of all finite subsets of $V(X)$. For each $a\in I$, define $I_{a}:=\{b\in I:a\subseteq b\}$. The family $\{I_{a}:a\in I\}$ has the f.i.p., so there is therefore an ultrafilter $\mathcal{U}$ over $I$ such that each $I_{a}\in\mathcal{U}$. Let $Y:=X^{I}/\mathcal{U}$. The ultrapower construction $V(X)\overset{\iota}{\hookrightarrow}W\overset{\cdot/\mathcal{U}}{\to}V(Y)$ given in the proof of Theorem 6.1 is an enlargement.
@@
and hence $I_{\{B\}}\subseteq\{a\in I:f(a)\in B\}$. It follows that $f/\mathcal{U}\in c_{B}/\mathcal{U}={}^{*}B$. As $B\in\mathcal{B}$ is arbitrary, it follows that $f/\mathcal{U}\in\bigcap_{B\in\mathcal{B}}{}^{*}B=\bigcap{}^{\sigma}\mathcal{B}$. As $\mathcal{B}$ is arbitrary, it follows that $\bigcap{}^{\sigma}\mathcal{B}\neq\varnothing$ for all $\mathcal{B}\in V(X)$ with the f.i.p, and thus by Theorem 5.3, the extension $V(X)\to V(Y)$ is an enlargement.
@@ 6.3 Existence of Comprehensive Transfer Maps
Let $V(X)$ be a superstructure over $X$, and let $V(X)\to V({}^{*}X)$ be the transfer map provided by an ultrapower construction, as in Section 6.1.1. Then * is comprehensive.
@@ 6.4 Existence of Polysaturated Extensions via Ultrapowers
We first show that the following claim holds:
Claim: If $\mathcal{L}$ is a countable language, and if $\Sigma(x)$ is a set of $\mathcal{L}$-formulas which is finitely satisfiable in an ultraproduct $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$, then $\Sigma(x)$ is satisfiable.
So suppose that $\Sigma(x)$ is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Since $\mathcal{L}$ is countable, so is $\Sigma(x)$, and hence we can enumerate it:
@@
[…]
@@
Observe that if $\mathcal{U}$ satisfies (S) then any ultraproduct modulo $\mathcal{U}$ interpreting a language $\mathcal{L}$ of cardinality $\kappa$ is $\kappa$-saturated. To see this, suppose that $X\subseteq\prod_{I}A_{i}/\mathcal{U}$ has $|X|<\kappa$. Then $\mathcal{L}_{X}$ is also a language of cardinality $\kappa$. Automatically, therefore, any set of $\mathcal{L}_{X}$-formulas $\Sigma(x)$ has $|\Sigma(x)|<\kappa$. Condition (S) then immediately yields that any set of $\mathcal{L}_{X}$-formulas $\Sigma(x)$ which is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$.
@@
Thus let $\Sigma(x)$ be a set of formulas of cardinality $\kappa$ which is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Then if $\Theta$ is a finite subset of $\Sigma$, the set $\{i\in I:\mathfrak{A}_{i}\vDash\exists x(\bigwedge\Theta)\}$ belongs to $\mathcal{U}$, by Łos’ Theorem. We thus have a map
@@
We say that a sequence $(\Phi_{i})_{i\in I}$ of members of $\mathcal{P}^{<\omega}(\Sigma)$ supports $p$ if it satisfies (i), (ii) of Lemma 6.11. Thus if there is a sequence supporting the relation $p$ defined in $(\star)$ above, then $\Sigma$ is satisfiable in $\prod_{i\in I}\mathfrak{A}_{i}/\mathcal{U}$.
@@
Observe that the order-reversal defined in $(\star)$ need not be either anti-additive or locally finite. However, if an abstract order-reversal $p$ has a support $\Phi:=(\Phi_{i})_{i\in I}$, then we can define another order-reversal $p_{\Phi}$ which has those properties:
@@
We now seek conditions that will ensure that any order-reversal on any set $X$ of cardinality $\kappa$ has a support. Observe that if $p,p^{\prime}$ are order-reversals so that $p^{\prime}\leq p$, then $p^{\prime}$ is locally finite if $p$ is, since $\{\Theta:i\in p^{\prime}(\Theta)\}\subseteq\{\Theta:i\in p(\Theta)\}$ for all $i\in I$. Furthermore, if $\Phi$ is a support of $p^{\prime}$, then it is a support of $p$ also, since $I=\{i:i\in p^{\prime}(\Phi_{i})\}\subseteq\{i:i\in p(\Phi_{i})\}$. Hence if we can define for every order-reversal $p$ two smaller order -eversals $Lp\leq p,Cp\leq p$ so that $Lp$ is locally finite, and $Cp$ is anti-additive, then $CLp\leq Lp\leq p$ is anti-additive and locally finite. Hence $CLp$ has a support, and this will be a support of $p$ also.
@@
  • Suppose that $\mathcal{U}$ is a countably incomplete $\kappa$-good ultrafilter on a set $I$, and that $(\mathfrak{A}_{i})_{i\in I}$ is a family of models interpreting a language $\mathcal{L}$ of cardinality $\kappa$.
@@
If $\mathcal{U}$ is a countably incomplete $\kappa$-good ultrafilter, then every ultraproduct modulo $\mathcal{U}$ interpreting a language of cardinality $\kappa$ is $\kappa$-saturated.